Six years ago Olwethu lent Happy R150 000 on condition that he would pay her back in nine years time. The applicable interest rate is 15,5% per year, compounded monthly. Happy also owes Olwethu another amount of R250 000 that he has to pay back six years from now for a loan that earned interest at 16,4% per year, compounded semi-annually. Happy asks Olwethu if he an settle both his debts three years from now. The total amount that Happy will have to pay Olwethu three years from now is


1
Expert's answer
2022-03-16T17:47:37-0400

If Happy borrows R150000 six years ago and to repay in three years time, and also borrows R250000 now, to also repay in six years time.

The amount he was supposed to pay back on the first R150000 is:

"FV=PV (1+\frac{r}{m})^{nm}"

Where n is 9

m is 12

"FV=R150000(1+\frac{
15.5\%}{12})^{9×12}"

"FV=R599863.88"


On the second R250000

"FV=R250000(1+\frac{16.4\%}{2})^{2×3}"

"FV=R401147.02"

The total amount he'd repay Olwethu three years from now is:

"=R599863.88+R401147.02"

"=R1001010.90"